OpenStudy (gabylovesyou):

Choose the correct simplification of the expression (−2x + 6y)(4x − 6y).

OpenStudy (gabylovesyou):

@thearcane

OpenStudy (johnweldon1993):

well use foil first. (-2x + 6y)(4x - 6y) means -2x* 4x -2x * -6y 6y * 4x 6y * -6y what do each of those equal?

OpenStudy (gabylovesyou):

-2x * 4x = -8x -2x * -6y = 12x 6y *4x = 24x 6y *-6y = 36y

OpenStudy (johnweldon1993):

remember when you multiply 2 x's...you get x * x...which is x²

OpenStudy (johnweldon1993):

when you multiply an x and a y...you just bring both terms over...since they cannot be combined

OpenStudy (gabylovesyou):

ok

OpenStudy (johnweldon1993):

so -2x * 4x = -8x² -2x * -6y = 12xy 6y *4x = 24xy 6y *-6y = -36y² right?

OpenStudy (johnweldon1993):

so if we have those as our answers.... -8x² + 12xy + 24xy -36y² what can be combined here? can x² be combined with xy? No...only another x² can xy be combined with y²? No just another xy but look...we have another xy in our equation....so those 2 can be combined. so what is 12xy + 24xy?

OpenStudy (gabylovesyou):

36xy

OpenStudy (gabylovesyou):

@johnweldon1993

OpenStudy (johnweldon1993):

correct... so alltogether we have -8x² + 36xy - 36y² is that one of the choices?

OpenStudy (gabylovesyou):

yess :) thank you

OpenStudy (johnweldon1993):

anytime :) glad to help :)